# Optimized 3D Gaussian Splatting using Coarse-to-Fine Image Frequency Modulation

### v1

@@ Abstract
The field of Novel View Synthesis has been revolutionized by 3D Gaussian Splatting (3DGS), which enables high-quality scene reconstruction that can be rendered in real-time. 3DGS-based techniques typically suffer from high GPU memory and disk storage requirements which limits their practical application on consumer-grade devices. We propose Opti3DGS, a novel frequency-modulated coarse-to-fine optimization framework that aims to minimize the number of Gaussian primitives used to represent a scene, thus reducing memory and storage demands. Opti3DGS  leverages image frequency modulation, initially enforcing a coarse scene representation and progressively refining it by modulating frequency details in the training images. On the baseline 3DGS [9], we demonstrate an average reduction of 62% in Gaussians, a 40% reduction in the training GPU memory requirements and a 20% reduction in optimization time without sacrificing the visual quality. Furthermore, we show that our method integrates seamlessly with many 3DGS-based techniques [19, 12, 5, 4, 15], consistently reducing the number of Gaussian primitives while maintaining, and often improving, visual quality. Additionally, Opti3DGS inherently produces a level-of-detail scene representation at no extra cost, a natural byproduct of the optimization pipeline. Results and code will be made publicly available.
@@
Figure 1: Opti3DGS progressively grows levels to reduce redundant and unused Gaussians while maintaining accurate geometric structure and competitive quality.
1 Introduction
@@
Photorealistic scene reconstruction remains a fundamental challenge in 3D computer vision, with applications spanning augmented and virtual reality, media production, and immersive gaming. Despite significant advancements, existing Novel View Synthesis (NVS) methods - particularly those based on Neural Radiance Fields (NeRF) [16] - face substantial computational constraints [8, 13, 2]. These constraints hinder broader adoption, especially on web, mobile, and edge devices, where efficient rendering and low memory usage are critical.
@@
3D Gaussian Splatting (3DGS) [9], recently introduced as an explicit point-based NVS technique, achieves high visual fidelity with fast training and real-time rendering. However, it suffers from significant GPU memory and storage overhead [19, 18], limiting its applicability across diverse device form factors.
@@
To address these challenges we introduce Opti3DGS, a novel and efficient coarse-to-fine optimization strategy that reduces the number of Gaussians used to represent a scene. This translates into reduced memory and storage overheads compared to 3DGS [9]. Our approach begins with intentional over-reconstruction, which serves as an effective regularizer, followed by progressive refinement of the scene to balance detail and efficiency. Several techniques have been proposed which aim to reduce 3DGS memory and storage costs. Some focus on compressing the stored representation [18, 12, 3, 17, 14], others optimize training speed and parallelization [22], while a few attempt to minimize Gaussian primitive count during optimization itself [12, 19]. Our approach falls into the latter category, reducing Gaussian primitives during the optimization process, thereby decreasing GPU memory demands and accelerating training without sacrificing visual quality. Notably, our method achieves these benefits without introducing additional learnable parameters or requiring lengthy optimizations.
@@
Opti3DGS improves the baseline 3DGS [9], achieving a 62% reduction in Gaussian primitives, a 40% decrease in GPU memory usage during training, and a 20% speedup in optimization time. Our technique is highly modular and can be easily incorporated into most 3DGS-based pipelines with minimal modifications. Furthermore, we demonstrate its positive impact when integrated into advanced 3DGS methods such as MiniSplatting [4] and Taming3DGS [15], as well as other recent appraches [19, 12, 5]. Due to the coarse-to-fine nature of our approach, it naturally produces level-of-detail representation as a byproduct, shown in Figure 1. To summarize, our key contributions are:
@@
Figure 2: Method Overview: We start the optimization with very large and coarse Gaussians which are then densified based on the differentiable rendering signal. Specifically we control the frequency spectrum of the training images, gradually increasing the threshold for the low pass filtering.
2 Related Work
@@
3DGS Compression: One branch of approaches that aim to reduce the storage size of 3DGS scenes focuses on quantizing raw parameters stored for each Gaussian primitive [18, 12, 3, 17, 14]. These enable efficient storage and transmission but since the optimization is done after 3DGS scene optimization, it does not lower GPU memory requirements during training and some require cumbersome decoding in GPU memory from a compressed representation which can be computationally expensive to render [10]. Some methods like [5] also include an efficient encoding and decoding mechanism for parameters during the optimization process, leading to both storage and compute time gains. However, in the case of [5] they use a separate neural network to decode the color values and face complications due to quantized vectors not being differentiable. A subset of these methods [18, 4, 20, 19] also compute the rendering importance of each Gaussian primitive based on its contribution to individual pixels, its relative size or the Gaussian neighborhood it is located in.
@@
Gaussian Primitive Reduction: An alternative approach is to actively reduce the number of primitives during the optimization process [12, 19, 4, 15]. The Adaptive Density Control (ADC) mechanism in [9] removes Gaussians below a fixed opacity threshold at predefined intervals and densifies ones with large gradients by either splitting or cloning them. But despite these measures the primitive count increases rapidly and with much variation across runs. [12] proposed to learn a volume aware mask during training, incurring additional learnable parameters and optimization time, to remove primitives which are either small or have minimal contribution. But this approach is limited in its effectiveness because the problem stems from dense clustering of Gaussians in certain locations with a significant number of them not even contributing to the color at all [18, 4]. [4] show that there are multi-view inconsistencies between the centroids of the Gaussians and their location density is not uniform even in areas with uniform color and texture attributes and propose to regularize this inconsistency. [19] prunes Gaussians during the optimization process based on their rendering importance and signs of local clustering, however they do not reduce the peak GPU memory requirements during the optimization process. [15] propose the use of per-view per-pixel saliency maps to guide the densification process in a selective manner with predictable growth trend and used fused operations to speed up optimization. However all of these approaches require major changes in the ADC functionality and thus are not easily inter-compatible with each other and have unknown effects if combined.
@@
Summary: We propose Opti3DGS, a coarse-to-fine frequency modulation approach to reduce the number of Gaussians by only changing the training images during the optimization process. [5] also use a coarse-to-fine strategy but the downsize images in resolution progressively, in contrast Opti3DGS modulates the image frequency. Both approaches have a distinct impact on the image pixels and quite different results. [21] also use a frequency-based approach for optimizing 3DGS but they provide gradient signals in multiple separate frequency bands simultaneously with the goal of increasing visual quality which leads to a significantly higher Gaussian count. Opti3DGS on the other hand sits behind the ADC and is not directly involved in the differential rendering process. This ensures wide ranging compatibility with most 3DGS based methods and allows Opti3DGS to integrate with compression and quantization-based methods. Opti3DGS also directly minimize the Gaussians which leads to reduced GPU memory requirements while the optimization process is still ongoing.
3 Method
Figure 3: We start with uniform distribution of larger Gaussians as compared to 3DGS which starts heavy densification from the very start of optimization thus complicating the loss landscape. The zoomed views show efficient use of the initial sparse pointcloud by our approach.
3.1 Preliminaries: 3D Gaussian Splatting
@@
3DGS [9] uses anisotropic Gaussians whose shape, color, opacity and refraction parameters are estimated via multi-view image-based optimization. Each Gaussian is defined by a covariance matrix $\Sigma$, a center position $\mathbf{p}$, opacity $\alpha$, and spherical harmonics coefficients for color $\mathbf{c}$. The covariance matrix $\Sigma$ is decomposed into a scaling matrix and a rotation matrix to facilitate differentiable optimization. For rendering, all Gaussians in the view frustum are splatted using onto a 2D plane followed by $\alpha$-blending of the color based on Gaussian opacity to achieve the final image. The color $C$ of a pixel is computed by blending $N$ ordered 2D Gaussians that overlap the pixel, formulated as:
$\footnotesize C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (1)
@@ 3.2 Opti3DGS Overview
3.3 Image Frequency Modulation
@@
3.4 Frequency Training Schedule
Starting with very large coarse Gaussians and progressively adding high frequency details requires a schedule of frequency removal and addition. We define various mechanisms, including linear, cosine, cosine restarting, exponential and step functions, as shown in Figure 4, which control the intensity of the image frequency modulation given the current training iteration as input. For the step function we define a single value for a fixed iteration interval and define all intervals accordingly. Based on ablation experiments, described in Section 4.3, the step function to be most effective.
Figure 4: Visualization of different decay rates reported in our work. The vertical axis shows the current kernel size for the filtering algorithm for any given iteration. Note we stop all frequency modulation after 12,000 iterations and pass full sized images for the remaining 18,000 iterations.
3.5 Progressive Growing Pipeline
Opti3DGS employs a progressive frequency control strategy with five distinct image and scene quality levels $I_{n}$ and $G_{n}$ respectively in range $n=(1-5)$ as depicted in Fig.1. We start by first obtaining a sparse scene reconstruction and initial 3D points using SfM on the training images $I$ and this initial sparse representation acts as our $G_{0}$ at $n=0$ for initializing the Gaussians.
$G_{0}=SfM(I)$ (3)
Then for each subsequent level in our coarse-to-fine optimization process we use the last set of Gaussians $G_{n-1}$ as the starting point and the normal 3DGS optimization loss is used to optimize $G_{n-1}$ using the filtered images $I$ in range $n=(1-5)$.
$G_{n}=3DGS({G_{n-1}},I_{n})$ (4)
@@ 3.5 Progressive Growing Pipeline
$G_{5}=3DGS({G_{4}},I)$ (5)
For level-of-detail representation results, the model is saved at the last iteration before transitioning to the next level.
Table 1: Opti3DGS consistently lowers Gaussian count for all reported methods as well as improving the reconstruction quality for some. Note we also report the maximum Gaussian count for methods where it peaks before the end of optimization.
Method Mip-NeRF360[1] Tanks & Temples[11] Deep Blending[7]
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
Compact-3DGS[12] 0.799 27.01 0.242 16:53 1.489 2.605 0.834 23.34 0.199 10:33 0.839 1.472 0.904 29.79 0.254 15:41 1.052 2.300
Compact-3DGS + Ours 0.792 26.96 0.262 13:37 1.048 1.697 0.827 23.26 0.217 08:31 0.601 1.016 0.904 29.81 0.258 13:46 0.831 1.710
Reduced-3DGS[19] 0.810 27.29 0.229 11:55 1.460 - 0.841 23.52 0.186 06:45 0.568 - 0.903 29.63 0.248 11:00 0.925 -
Reduced-3DGS + Ours 0.801 27.14 0.250 09:55 1.061 - 0.834 23.52 0.203 05:38 0.409 - 0.904 29.76 0.251 09:53 0.697 -
Mini-Splatting[4] 0.820 27.25 0.219 11:01 0.496 4.240 0.832 23.22 0.202 07:10 0.203 4.311 0.904 29.88 0.256 10:00 0.352 4.517
Mini-Splatting + Ours 0.821 27.26 0.219 11:21 0.488 4.209 0.834 23.17 0.204 07:13 0.197 4.223 0.908 30.01 0.254 10:07 0.344 4.387
Taming3DGS[15] (budget) 0.793 27.21 0.262 14:36 0.666 - 0.832 23.66 0.213 02:19 0.319 - 0.899 29.75 0.273 02:14 0.294 -
Taming3DGS (budget) + Ours 0.788 27.27 0.274 14:11 0.5735 - 0.834 23.90 0.215 02:17 0.318 - 0.905 30.00 0.269 02:22 0.294 -
Taming3DGS[15] (big) 0.823 27.82 0.207 06:46 3.251 - 0.859 24.14 0.160 05:04 2.290 - 0.909 30.04 0.233 06:15 3.593 -
Taming3DGS (big) + Ours 0.817 27.77 0.223 05:19 2.345 - 0.854 24.22 0.175 03:38 1.333 - 0.911 30.13 0.237 04:57 2.212 -
EAGLES[5] 0.81 27.23 0.24 09:40 1.33 - 0.84 23.27 0.20 06:19 0.65 - 0.91 29.86 0.25 11:22 1.19 -
EAGLES + Ours 0.79 26.90 0.27 13:21 0.90 - 0.83 23.28 0.21 07:00 0.58 - 0.91 29.86 0.25 11:36 0.87 -
3DGS[9] 0.813 27.53 0.221 11:42 3.157 - 0.844 23.69 0.178 07:19 1.825 - 0.900 29.54 0.247 11:34 2.795 -
3DGS + Ours 0.809 27.09 0.237 09:50 1.327 - 0.823 23.15 0.217 05:56 0.658 - 0.902 29.56 0.255 10:26 0.953 -
4 Results and Evaluation
Datasets: Following common practice we use the MipNeRF360 [1], DeepBlending [7] and Tanks and Temples [11] datasets with a total of 13 scenes including a good mix of indoor and outdoor unbounded scenes. We use the train and truck scene from [11], the drjohnson and playroom scene from [7] and bicycle, bonsai, treehill, counter, kitchen, stump, flowers, garden and room scenes from [1].

### v2

@@ Abstract.
The field of Novel View Synthesis has been revolutionized by 3D Gaussian Splatting (3DGS), which enables high-quality scene reconstruction that can be rendered in real-time. 3DGS-based techniques typically suffer from high GPU memory and disk storage requirements which limits their practical application on consumer-grade devices. We propose Opti3DGS, a novel frequency-modulated coarse-to-fine optimization framework that aims to minimize the number of Gaussian primitives used to represent a scene, thus reducing memory and storage demands. Opti3DGS  leverages image frequency modulation, initially enforcing a coarse scene representation and progressively refining it by modulating frequency details in the training images. On the baseline 3DGS (Kerbl et al., 2023), we demonstrate an average reduction of 62% in Gaussians, a 40% reduction in the training GPU memory requirements and a 20% reduction in optimization time without sacrificing the visual quality. Furthermore, we show that our method integrates seamlessly with many 3DGS-based techniques (Papantonakis et al., 2024; Lee et al., 2024; Girish et al., 2024; Fang and Wang, 2024; Mallick et al., 2024), consistently reducing the number of Gaussian primitives while maintaining, and often improving, visual quality.
3D Gaussian Splatting, Efficient 3D Gaussian Splatting, Novel View Synthesis
journalyear: 2025copyright: acmlicensedconference: European Conference on Visual Media Production; December 3–4, 2025; London, United Kingdombooktitle: European Conference on Visual Media Production (CVMP ’25), December 3–4, 2025, London, United Kingdomdoi: 10.1145/3756863.3769707isbn: 979-8-4007-2117-5/2025/12ccs: Computing methodologies Computer graphics; Reconstruction
@@
Figure 1. Opti3DGS progressively grows levels to reduce redundant and unused Gaussians while maintaining accurate geometric structure and competitive quality.
1. Introduction
Figure 2. We start with uniform distribution of larger Gaussians as compared to 3DGS which starts heavy densification from the very start of optimization thus complicating the loss landscape. The zoomed views show efficient use of the initial sparse pointcloud by our approach.
@@ 1. Introduction
Photorealistic scene reconstruction remains a fundamental challenge in 3D computer vision, with applications spanning augmented and virtual reality, media production, and immersive gaming. Despite significant advancements, existing Novel View Synthesis (NVS) methods - particularly those based on Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) - face substantial computational constraints (Hu et al., 2022; Li et al., 2023; Deng and Tartaglione, 2023). These constraints hinder broader adoption, especially on web, mobile, and edge devices, where efficient rendering and low memory usage are critical.
@@
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), recently introduced as an explicit point-based NVS technique, achieves high visual fidelity with fast training and real-time rendering. However, it suffers from significant GPU memory and storage overhead (Papantonakis et al., 2024; Niedermayr et al., 2024), limiting its applicability across diverse device form factors.
@@
To address these challenges we introduce Opti3DGS, a novel and efficient coarse-to-fine optimization strategy that reduces the number of Gaussians used to represent a scene. This translates into reduced memory and storage overheads compared to 3DGS (Kerbl et al., 2023). Our approach begins with intentional over-reconstruction, which serves as an effective regularizer, followed by progressive refinement of the scene to balance detail and efficiency. Several techniques have been proposed which aim to reduce 3DGS memory and storage costs. Some focus on compressing the stored representation (Niedermayr et al., 2024; Lee et al., 2024; Fan et al., 2024; Navaneet et al., 2023; Liu et al., 2024), others optimize training speed and parallelization (Zhao et al., 2025), while a few attempt to minimize Gaussian primitive count during optimization itself (Lee et al., 2024; Papantonakis et al., 2024). Our approach falls into the latter category, reducing Gaussian primitives during the optimization process, thereby decreasing GPU memory demands and accelerating training speed while maintaining visual quality. Notably, our method achieves these benefits as a plug-in approach without introducing additional learnable parameters or requiring lengthy optimizations.
@@
Opti3DGS improves the baseline 3DGS (Kerbl et al., 2023), achieving a 62% reduction in Gaussian primitives, a 40% decrease in GPU memory usage during training, and a 20% speedup in optimization time. Our technique is highly modular and can be easily incorporated into most 3DGS-based pipelines with minimal modifications. Furthermore, we demonstrate its positive impact when integrated into optimized 3DGS variants such as MiniSplatting (Fang and Wang, 2024) and Taming3DGS (Mallick et al., 2024), as well as other recent appraches (Papantonakis et al., 2024; Lee et al., 2024; Girish et al., 2024). Due to the coarse-to-fine nature of our approach, it naturally produces level-of-detail representation as a byproduct, shown in Figure 1, which could have useful applications in foveated rendering, streaming applications and advanced storage mechanisms. To summarize, our key contributions are:
@@
2. Related Work
@@
3DGS Compression: One approach to reduce the storage size of 3DGS scenes focuses on quantizing raw parameters stored for each Gaussian primitive (Niedermayr et al., 2024; Lee et al., 2024; Fan et al., 2024; Navaneet et al., 2023; Liu et al., 2024). These enable efficient storage and transmission but since the optimization is done after 3DGS scene optimization, it does not lower GPU memory requirements during training and some of these methods require cumbersome decoding in GPU memory from a compressed representation which can be computationally expensive to render (Kim et al., 2024). Some methods like (Girish et al., 2024) also include an efficient encoding and decoding mechanism for parameters during the optimization process, leading to both storage and compute time gains. However, in the case of (Girish et al., 2024) they introduce additional learnable parameters and use a separate neural network to decode the color values and face complications due to quantized vectors not being differentiable. (Girish et al., 2024) also use progressive resolution training which is different from our frequency based image filtering as outlined in section 4. A subset of these methods (Niedermayr et al., 2024; Girish et al., 2024; Fang and Wang, 2024; Wang et al., 2024; Papantonakis et al., 2024) also compute the rendering importance of each Gaussian primitive based on its contribution to individual pixels, its relative size or the Gaussian neighborhood it is located in.
@@
Gaussian Primitive Reduction: An alternative approach is to actively reduce the number of primitives during the optimization process (Lee et al., 2024; Papantonakis et al., 2024; Fang and Wang, 2024; Mallick et al., 2024). The Adaptive Density Control (ADC) mechanism in (Kerbl et al., 2023) removes Gaussians below a fixed opacity threshold at predefined intervals and densifies ones with large gradients by either splitting or cloning them. But despite these measures the primitive count increases rapidly and with much variation across runs(Mallick et al., 2024). (Lee et al., 2024) proposed to learn a volume aware mask during training, incurring additional learnable parameters and optimization time, to remove primitives which are either small or have minimal contribution. But this approach is limited in its effectiveness because the problem stems from dense clustering of Gaussians in certain locations with a significant number of them not even contributing to the color at all (Niedermayr et al., 2024; Fang and Wang, 2024). (Fang and Wang, 2024; Yang et al., 2025) show that there are multi-view inconsistencies between the centroids of the Gaussians and their location density is not uniform even in areas with uniform color and texture attributes and propose to regularize this inconsistency. (Papantonakis et al., 2024) prunes Gaussians during the optimization process based on their rendering importance and signs of local clustering, however they do not reduce the peak GPU memory requirements during the optimization process. (Mallick et al., 2024) propose the use of per-view per-pixel saliency maps to guide the densification process in a selective manner with predictable growth trend and used fused operations to speed up optimization. However all of these approaches require major changes in the ADC functionality and thus are not easily inter-compatible with each other and have unknown effects if combined. Whereas, Opti3DGS  provides a plug-in approach that can be combined with various pipelines with almost no change to the core algorithm.
@@
(Girish et al., 2024) also use a coarse-to-fine strategy but the downsize images in resolution progressively, in contrast Opti3DGS modulates the image frequency with the image resolution staying fixed. Both approaches have a distinct impact on the image pixels and quite different results. (Zhang et al., 2024) also use a frequency-based approach for optimizing 3DGS but they provide gradient signals in multiple separate frequency bands simultaneously with the goal of increasing visual quality which leads to a significantly higher Gaussian count.
Summary: We propose Opti3DGS, a coarse-to-fine frequency modulation approach to reduce the number of Gaussians by only changing the training images during the optimization process. Opti3DGS  sits behind the ADC and is not directly involved in the differential rendering process and no changes need to be made in the core algorithm. This ensures wide ranging plug-in compatibility with most 3DGS based methods and also allows Opti3DGS to integrate with compression and quantization-based methods. Opti3DGS also directly minimizes the Gaussians which leads to reduced GPU memory requirements during the optimization process reducing overall GPU memory requirements.
Figure 3. Method Overview: We start the optimization with very large and coarse Gaussians which are then densified based on the differentiable rendering signal. Specifically we control the frequency spectrum of the training images, gradually increasing the threshold for the low pass filtering.
3. Method
3.1. Preliminaries: 3D Gaussian Splatting
@@
3DGS (Kerbl et al., 2023) uses anisotropic Gaussians whose shape, color, opacity and refraction parameters are estimated via multi-view image-based optimization. Each Gaussian is defined by a covariance matrix $\Sigma$, a center position $\mathbf{p}$, opacity $\alpha$, and spherical harmonics coefficients for color $\mathbf{c}$. The covariance matrix $\Sigma$ is decomposed into a scaling matrix and a rotation matrix to facilitate differentiable optimization. For rendering, all Gaussians in the view frustum are splatted using onto a 2D plane followed by $\alpha$-blending of the color based on Gaussian opacity to achieve the final image. The color $C$ of a pixel is computed by blending $N$ ordered 2D Gaussians that overlap the pixel, formulated as:
(1) $\footnotesize C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$
@@ 3.2. Opti3DGS Overview
3.3. Image Frequency Modulation
@@
3.4. Frequency Training Schedule
Starting with very large coarse Gaussians and progressively adding high frequency details requires a schedule of frequency removal and addition. We explore various mechanisms, including linear, cosine, cosine restarting, exponential and step functions, as shown in Figure 4, which control the intensity of the image frequency modulation given the current training iteration as input. For the step function we define a single value for a fixed iteration interval and define all intervals accordingly. Based on ablation experiments, described in Section 4.3, the step function to be most effective.
Figure 4. Visualization of different decay rates reported in our work. The vertical axis shows the current kernel size for the filtering algorithm for any given iteration. Note we stop all frequency modulation after 12,000 iterations and pass full sized images for the remaining 18,000 iterations.
3.5. Progressive Growing Pipeline
Opti3DGS employs a progressive frequency control strategy with five distinct image and scene quality levels $I_{n}$ and $G_{n}$ respectively in range n=(15)n=(1-5) as depicted in Fig.1. We start by first obtaining a sparse scene reconstruction and initial 3D points using SfM on the training images $I$ and this initial sparse representation acts as our $G_{0}$ at $n=0$ for initializing the Gaussians.
(3) $G_{0}=SfM(I)$
Then for each subsequent level in our coarse-to-fine optimization process we use the last set of Gaussians $G_{n-1}$ as the starting point and the normal 3DGS optimization loss is used to optimize $G_{n-1}$ using the filtered images $I$ in range n=(15)n=(1-5).
(4) $G_{n}=3DGS({G_{n-1}},I_{n})$
@@ 3.5. Progressive Growing Pipeline
(5) $G_{5}=3DGS({G_{4}},I)$
For level-of-detail representation results, the model for any given level is saved at the last iteration before transitioning to the next level. The number of levels as well as the decay schedule can be easily changed for downstream applications.
Figure 5. We reduce the GPU memory required during optimization due to a reduced Gaussian primitive load, which also reduces the total optimization time. These results are for 3DGS(Kerbl et al., 2023) across the bicycle scenes and are represntive of all scenes tested.
Figure 6. Pixel level comparison of error maps against the ground truth image. Our approach maintains similar quantitative scores while using fewer Gaussians and results in a decreased pixel wise error. Note the absence of noise like bright points in our results, due to the removal of tiny redundant Gaussians.
[…]
Method Deep Blending(Hedman et al., 2021) Tanks & Temples(Knapitsch et al., 2017) Mip-NeRF360(Barron et al., 2022)
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
@@
Compact-3DGS(Lee et al., 2024) 0.904 29.79 0.254 274 15:41 1.052 1.428 0.834 23.34 0.199 150 10:33 0.839 1.006 0.799 27.01 0.242 275 16:53 1.489 2.400
Compact-3DGS + Ours 0.904 29.81 0.258 348 13:46 0.831 (-21%) 1.130 0.827 23.26 0.217 218 08:31 0.601 (-28%) 0.718 0.792 26.96 0.262 357 13:37 1.048 (-30%) 1.794
Reduced-3DGS(Papantonakis et al., 2024) 0.903 29.63 0.248 420 11:00 0.925 - 0.841 23.52 0.186 258 06:45 0.568 - 0.810 27.29 0.229 408 11:55 1.460 -
Reduced-3DGS + Ours 0.904 29.76 0.251 558 09:53 0.697 (-25%) - 0.834 23.52 0.203 352 05:38 0.409 (-28%) - 0.801 27.14 0.250 560 09:55 1.061 (-27%) -
Mini-Splatting(Fang and Wang, 2024) 0.904 29.88 0.256 839 10:00 0.352 0.378 0.832 23.22 0.202 602 07:10 0.203 0.217 0.820 27.25 0.219 699 11:01 0.496 0.611
Mini-Splatting + Ours 0.908 30.01 0.254 849 10:07 0.344 (-2%) 0.376 0.834 23.17 0.204 614 07:13 0.197 (-3%) 0.211 0.821 27.26 0.219 711 11:21 0.488 (-2%) 0.610
Taming3DGS(Mallick et al., 2024) (budget) 0.899 29.75 0.273 1038 02:14 0.294 - 0.832 23.66 0.213 411 02:19 0.319 - 0.793 27.21 0.262 523 14:36 0.666 -
Taming3DGS (budget) + Ours 0.905 30.00 0.269 1036 02:22 0.294 (0%) - 0.834 23.90 0.215 431 02:17 0.318 (0%) - 0.788 27.27 0.274 593 14:11 0.5735 (-14%) -
Taming3DGS(Mallick et al., 2024) (big) 0.909 30.04 0.233 326 06:15 3.593 - 0.859 24.14 0.160 107 05:04 2.290 - 0.823 27.82 0.207 228 06:46 3.251 -
Taming3DGS (big) + Ours 0.911 30.13 0.237 431 04:57 2.212 (-38%) - 0.854 24.22 0.175 163 03:38 1.333 (-42%) - 0.817 27.77 0.223 318 05:19 2.345 (-28%) -
EAGLES(Girish et al., 2024) 0.91 29.86 0.25 247 11:22 1.19 - 0.84 23.27 0.20 379 06:19 0.65 - 0.81 27.23 0.24 274 09:40 1.33 -
EAGLES + Ours 0.91 29.86 0.25 304 11:36 0.87 (-27%) - 0.83 23.28 0.21 417 07:00 0.58 (-11%) - 0.79 26.90 0.27 329 13:21 0.90 (-32%) -
3DGS(Kerbl et al., 2023) 0.900 29.54 0.247 293 11:34 2.795 - 0.844 23.69 0.178 108 07:19 1.825 - 0.813 27.53 0.221 247 11:42 3.157 -
3DGS + Ours 0.902 29.56 0.255 611 10:26 0.953 (-66%) - 0.823 23.15 0.217 258 05:56 0.658 (-64%) - 0.809 27.09 0.237 308 09:50 1.327 (-58%) -
4. Results and Evaluation
Datasets: We report results on the MipNeRF360 (Barron et al., 2022), DeepBlending (Hedman et al., 2021) and Tanks and Temples (Knapitsch et al., 2017) datasets with a total of 13 scenes which include a good mix of indoor and outdoor unbounded scenes, as well as scenes with low-frequency textures.